# Deciphering interventional dynamical causality from non-intervention systems

### v1

@@
Deciphering interventional dynamical causality from non-intervention systems
Jifan Shi1,2,3, Yang Li4, Juan Zhao5, Siyang Leng1,6, Kazuyuki Aihara4∗, Luonan Chen4,7,8∗, Wei Lin1,2,3∗
1Research Institute of Intelligent Complex Systems, Fudan University, Shanghai 200433, China
2Shanghai Artificial Intelligence Laboratory, Shanghai 200232, China
3State Key Laboratory of Medical Neurobiology and MOE Frontiers Center for Brain Science, Fudan University, Shanghai 200032, China
4International Research Center for Neurointelligence, The University of Tokyo Institutes for Advanced Study, The University of Tokyo,Tokyo 113-0033, Japan
5School of Pharmacy, Shanghai University of Traditional Chinese Medicine, Shanghai 201203, China
6Academy for Engineering and Technology, Fudan University, Shanghai 200433, China
7Key Laboratory of Systems Biology, Shanghai Institute of Biochemistry and Cell Biology, Center for Excellence in Molecular Cell Science, Chinese Academy of Sciences, Shanghai 200031, China
8Key Laboratory of Systems Health Science of Zhejiang Province, School of Life Science, Hangzhou Institute for Advanced Study, University of Chinese Academy of Sciences, Chinese Academy of Sciences, Hangzhou 310024, China
To whom correspondence should be addressed; E-mail: kaihara@g.ecc.u-tokyo.ac.jp or lnchen@sibs.ac.cn or wlin@fudan.edu.cn.
Detecting and quantifying causality is a focal topic in the fields of science, engineering, and interdisciplinary studies. However, causal studies on non-intervention systems attract much attention but remain extremely challenging. To address this challenge, we propose a framework named Interventional Dynamical Causality (IntDC) for such non-intervention systems, along with its computational criterion, Interventional Embedding Entropy (IEE), to quantify causality. The IEE criterion theoretically and numerically enables the deciphering of IntDC solely from observational (non-interventional) time-series data, without requiring any knowledge of dynamical models or real interventions in the considered system. Demonstrations of performance showed the accuracy and robustness of IEE on benchmark simulated systems as well as real-world systems, including the neural connectomes of C. elegans, COVID-19 transmission networks in Japan, and regulatory networks surrounding key circadian genes.
Introduction
@@
Statisticians believe that the causality is contained and obtainable in universal random variables regardless of the time label. The Neyman-Rubin’s potential outcome framework  (?, ?, ?), the Wright’s structural equation model  (?, ?) and the Pearl’s causal diagram model  (?, ?) are the most famous statistical approaches, which have been proved to be of mathematical equivalence  (?). Especially, the instrumental variable method, that was awarded the 2021 Nobel Prize in economic sciences, is widely applied in the study of causality among economics, environments, and other disciplines  (?). Statistical methods seek to discern binary causal relations among random variables on a directed acyclic graph, without relying on temporal data. However, in general complex dynamical systems, the causation must precede the effect, feedback is common, and the causal strength needs to be quantitatively measured. Time should play definitely a key role.
@@
Dynamics-oriented researchers have proposed fruitful algorithms for measuring the observational causality from time series, commonly referred to as the dynamical causality  (?). The celebrated Granger causality (GC)  (?, ?) employs a linear model and uses improvement of predictability over time to illustrate the causality. Transfer entropy (TE)  (?) generalizes GC to the nonlinear case by quantifying the prediction uncertainty through Shannon entropy. Neither GC nor TE addresses the “non-separability” problem, which means that removing the causal variable from the system inevitably influences the dynamics of downstream variables  (?, ?, ?, ?). To measure causality in non-separable systems, numerous approaches have emerged in the last decade within the framework of delay embedding. These include convergence of cross mapping (CCM)  (?, ?), partial cross mapping (PCM)  (?), continuity scaling  (?), inverse continuity  (?), topological expansion  (?), joint distance distribution  (?), embedding entropy (EE)  (?) and other indices  (?, ?).
@@
Nonetheless, most indices designed for quantifying the dynamical causality primarily estimate directional causal relations at the observational level, i.e. observational dynamical causality (ObsDC). Conventionally, to detect and quantify causality at the interventional level, which is noted as the interventional dynamical causality (IntDC) in this study, intelligent modulation and manipulation of the dynamical system should be taken into account. By allowing external intervention to the system and recording data under different perturbations, frameworks such as perturbation cascade inference (PCI)  (?) and dynamical causal effect (DCE)  (?, ?) provide relevant computational schemes. However, due to ethical or practical limitations, many real systems should be analyzed without any external intervention. An essential problem is how to measure IntDC solely from the observational data but at the interventional level.
@@
In this study, we propose the IntDC framework and introduce a criterion named Interventional Embedding Entropy (IEE), which aims to identify and quantify IntDC between variables solely from the observational time series. Actually, IEE is rigorously derived by the theory of the delay embedding. Numerical IEE does not require specific prior knowledge of dynamics, and additional perturbation to the system is also unnecessary. Compared to ObsDC indices, IEE designed for IntDC has the capability to rank the importance of causal effects and construct directional causal networks more effectively. We demonstrate numerical experiments on both simulated examples and real datasets, including estimating neural connectomes of C. elegans, evaluating COVID-19 transmission in Japan, and constructing regulatory networks surrounding key circadian genes.
Methods
Observational dynamical causality in delay-embedding space
We consider a discrete two-variable dynamical system as
@@ Methods
where $x,y$ are two variables, $p$ denotes the time step during which causality is considered, and $\varepsilon_{\cdot,t}$ stand for small noise terms. According to Shi et al.  (?), there exists ObsDC from $x$ to $y$ since the evolving equation of $y_{t+1}$ depends on the historical behavior of $x$, while there is no ObsDC from $y$ to $x$ as the dynamics of $x_{t+1}$ is independent of $y$. GC and TE measure the predictability of $f$ in the original time-series space to quantify the ObsDC from $x$ to $y$ through linear regression and entropy uncertainty, respectively (Fig. 1(B)). The causal strength of GC/TE is based on the assumption of separability, which means that removing $x$ from the system does not influence the observed data of $y$. However, in general coupled systems, the separability is not satisfied, as removing one variable changes the values of affected ones  (?, ?, ?).
To address the challenge of detecting ObsDC in universal non-separable systems, various approaches from the delay-embedding space have been proposed. In the delay-embedding framework, the bivariate system Eq. (1) is assumed to evolve into an attractive manifold with an inner dimension $d$. Let the time-delay vectors of $x$ and $y$ be
@@
where $L$ is the time-delay length, $\mathcal{M}_{X}$ and $\mathcal{M}_{Y}$ represent the manifolds formed by $\bm{X}_{t}$ and $\bm{Y}_{t}$, respectively. According to the seminal stochastic version of Takens’ embedding theorem  (?, ?, ?, ?), for delay dimension $L\geqslant 2d$, we can obtain
@@
where $\bm{F}$ is a smooth projection operator. The detailed derivation can be referred to the Supplementary Text. Equation (4) shows that the causal variable $\bm{X}_{t}\in\mathcal{M}_{X}$ can be reconstructed by the effect variable $\bm{Y}_{t+1}\in\mathcal{M}_{Y}$. However, as $\bm{F}$ is not reversible in generic sense, we can not determine $\bm{Y}_{t+1}$ only by the information from $\bm{X}_{t}$. The causal dependence in $f$ between $x$ and $y$ is transformed into the reconstructability of $\bm{F}$ (Fig. 1(D)). Instead of fitting a model by removing the causal variable as in GC and TE, detecting the existence and quantifying the continuity characteristics of $\bm{F}$ in the delay-embedding space can be adequate for causal identification, especially for universal non-separable systems. Related algorithms include CCM, PCM, and EE.
@@
Figures 1(A), (B) and (D) summarize dynamical causality at the observational level, where the ObsDC from variable $x$ to $y$ is estimated only from the historical observed time series. But fitting Eq. (1) or Eq. (4) from data is criticized as mere association or prediction  (?, ?).
@@
Figure 1: Illustration for the observational dynamical causality (ObsDC) and interventional dynamical causality (IntDC). (A) An illustrative example of a complex system where $(x_{t},y_{t})$ are observed time series. The $z_{t}$ stands for the unobserved variable. (B) ObsDC from $x$ to $y$ indicates that the dynamics of $y$ depends on $x$. Algorithms such as Granger causality (GC) and transfer entropy (TE) infer the ObsDC in the original time-series space. (C) IntDC from $x$ to $y$ indicates that the deviation of $y$, i.e. $\delta y$, depends on an intervention on $x$, i.e. $\delta x$. Usually additional data from a perturbed system $(\tilde{x}_{t},\tilde{y}_{t})$ is necessary for detecting IntDC. (D) ObsDC can be detected in the embedding space by cross mapping $\mathbb{F}$. The causal vector $X_{t}$ can be reconstructed by the effect vector $\bm{Y}_{t+1}$. (E) IntDC can be modeled in the embedding space. The interventional causal vector $\delta\bm{X}_{t}$ should be reconstructable by the effect vector $\delta\bm{Y}_{t+1}$ around $\bm{Y}_{t+1}$. The $\delta\bm{Y}_{t+1}$ is approximated by the distance from $\bm{Y}_{t+1}$ to its neighbors $\bm{Y}_{t_{k}+1}$, while $\delta\bm{X}_{t}$ is obtained by $X_{t_{k}}$ with the same time label $t_{k}$. (F) Interventional embedding entropy (IEE) measures the average information retention contained in $\delta\bm{Y}_{t+1}$ from $\delta\bm{X}_{t}$, quantifying the IntDC from variable $x$ to $y$. Notably, IEE does not require additional perturbations to the system and can be inferred from observational data solely.
Interventional dynamical causality in delay-embedding space
The detection of hidden essential-level causality necessitates intervention or manipulation to the system  (?, ?, ?, ?). We define that there is an IntDC from $x$ to $y$ during the time $[t-p,t]$ in the original time-series space, if $\delta y_{t}=\tilde{y}_{t}-y_{t}$ depends on $\delta x_{t}=\tilde{x}_{t}-x_{t}$, where the dynamics of $x_{t}$ is perturbed to $\tilde{x}_{t}$, and $\tilde{y}_{t}$ is the dynamics of affected $y$ after the perturbation (Fig. 1(C)). The intervention $\delta x_{t}$ can be either a pulse stimulation or persistent disturbances to the system.
To measure the IntDC, data before and after the intervention, i.e. $(x_{t},y_{t})$ and $(\tilde{x}_{t},\tilde{y}_{t})$, are typically required, as in PCI and DCE. These conventional algorithms rely on reproducible dynamics or data availability under different settings.
@@
To address challenges of causal inference in widespread non-intervention, non-linear and non-separable systems, we turn to the study of intervened dynamics Eq. (4) in the delay-embedding space. Denote $\delta\bm{X}_{t}=\widetilde{\bm{X}}_{t}-\bm{X}_{t}$ and $\delta\bm{Y}_{t+1}=\widetilde{\bm{Y}}_{t+1}-\bm{Y}_{t+1}$, where $\widetilde{\bm{X}}_{t}$ and $\widetilde{\bm{Y}}_{t+1}$ are the time-delay vectors of $x$ and $y$ after the intervention, respectively. Under infinitesimal intervention, Eq. (4) leads to
$\delta\bm{X}_{t}=\nabla_{\bm{Y}}\bm{F}(\bm{Y}_{t+1})\cdot\delta\bm{Y}_{t+1}.$ (5)
Equation (5) indicates that $\delta\bm{Y}_{t+1}$ contains complete information of $\delta\bm{X}_{t}$ in the neighborhood of $\bm{Y}_{t+1}$ or reconstructs $\delta\bm{X}_{t}$, if there exists IntDC from the causal variable $x$ to the effect variable $y$ (Fig. 1(E)).
Figures 1(A), (C), and (E) give an overview of the IntDC framework.
Interventional embedding entropy

### v2

# Deciphering Interventional Dynamical Causality from Non-intervention Complex Systems

@@
Deciphering Interventional Dynamical Causality from Non-intervention Complex Systems
Jifan Shi1∗, Yang Li2, Juan Zhao3, Siyang Leng1,4, Rui Bao5, Kazuyuki Aihara2∗, Luonan Chen6,7∗, Wei Lin1∗
1Research Institute of Intelligent Complex Systems & CISOR, Fudan University, Shanghai 200433, China
2International Research Center for Neurointelligence, The University of Tokyo Institutes for Advanced Study, The University of Tokyo, Tokyo 113-0033, Japan
3School of Pharmacy, Shanghai University of Traditional Chinese Medicine, Shanghai 201203, China
4Institute of AI and Robotics, College of Intelligent Robotics and Advanced Manufacturing, Fudan University, Shanghai 200433, China
5Frontiers Science Center for Deep Ocean Multispheres and Earth System, Key Laboratory of Marine Chemistry Theory and Technology, Ministry of Education, Ocean University of China, Qingdao 266100, China
6School of Mathematical Sciences and School of AI, Shanghai Jiao Tong University, Shanghai 200240, China
7Key Laboratory of Systems Health Science of Zhejiang Province, Hangzhou Institute for Advanced Study, University of Chinese Academy of Sciences, Chinese Academy of Sciences, Hangzhou 310024, China
To whom correspondence should be addressed. E-mail: jfshi@fudan.edu.cn; kaihara@g.ecc.u-tokyo.ac.jp; lnchen@sibs.ac.cn; wlin@fudan.edu.cn.
Detecting and quantifying causality is a focal topic in the fields of science, engineering, and interdisciplinary studies. However, causal studies on non-intervention systems attract much attention but remain extremely challenging. Delay-embedding technique provides a promising approach. In this study, we propose a framework named Interventional Dynamical Causality (IntDC) in contrast to the traditional Constructive Dynamical Causality (ConDC). ConDC, including Granger causality, transfer entropy and convergence of cross-mapping, measures the causality by constructing a dynamical model without considering interventions. A computational criterion, Interventional Embedding Entropy (IEE), is proposed to measure causal strengths in an interventional manner. IEE is an intervened causal information flow but in the delay-embedding space. Further, the IEE theoretically and numerically enables the deciphering of IntDC solely from observational (non-interventional) time-series data, without requiring any knowledge of dynamical models or real interventions in the considered system. In particular, IEE can be applied to rank causal effects according to their importance and construct causal networks from data. We conducted numerical experiments on Logistic dynamics, coupled-Henon maps, and chaotic neural networks to demonstrate that IEE can find causal edges accurately, eliminate effects of confounding, and quantify causal strength robustly over traditional indices. We also applied IEE to real-world tasks, including estimating neural connectomes of C. elegans, detecting COVID-19 transmission networks in Japan, and investigating regulatory networks surrounding key circadian genes. IEE performed as an accurate and robust tool for causal analyses solely from the observational data. The IntDC framework and IEE algorithm provide an efficient approach to the study of causality from time series in diverse non-intervention complex systems.
1 Introduction
@@
Statisticians believe that causality is contained and obtainable in universal random variables regardless of the time label. Neyman-Rubin’s potential outcome framework  (?, ?, ?), the Wright’s structural equation model  (?, ?) and the Pearl’s causal diagram model  (?, ?) are the most famous statistical approaches, which have been proved to be of mathematical equivalence  (?). Especially, the instrumental variable method is widely applied in the study of causality among economics, environments, and other disciplines  (?). Statistical methods seek to discern binary causal relations among random variables on a directed acyclic graph, without relying on temporal data. However, in general complex dynamical systems, the causation must precede the effect, feedback is common, and the causal strength needs to be quantitatively measured. Time should play definitely a key role.
@@
Dynamics-oriented researchers have proposed fruitful algorithms for measuring the observational causality from time series, commonly referred to as the dynamical causality  (?). The celebrated Granger causality (GC)  (?, ?) employs a linear model and uses improvement of predictability over time to illustrate the causality. Transfer entropy (TE)  (?) generalizes GC to the nonlinear case by quantifying the prediction uncertainty through Shannon entropy. Neither GC nor TE addresses the “non-separability” problem, which means that removing the causal variable from the system inevitably influences the dynamics of downstream variables  (?, ?, ?, ?, ?, ?). To measure causality in non-separable systems, numerous approaches have emerged in the last decade within the framework of delay embedding. These include convergence of cross mapping (CCM)  (?, ?), partial cross mapping (PCM)  (?), continuity scaling  (?), inverse continuity  (?), topological expansion  (?), joint distance distribution  (?), embedding entropy (EE)  (?) and other indices  (?, ?). Besides, Runge et al. designed the PCMCI method to further remove “common drivers” by selecting parent nodes iteratively (?). Friston et al. proposed the dynamical causal modeling (DCM), which is a Bayesian fitting from data with pre-selected models  (?). DCM was originally developed for modeling neural dynamics  (?). Krakovská et al. conducted a comparative analysis of six methods for detecting causality in bivariate systems (?).
@@
Nonetheless, most indices designed for quantifying causality from time series primarily focus on estimating directional causal relationships at the constructive level, referred to as constructive dynamical causality (ConDC). Traditionally, to detect and quantify causality at the interventional level, which is termed interventional dynamical causality (IntDC) in this study, requires intelligent modulation and manipulation of the dynamical system. By allowing external intervention to the system and recording data under different perturbations, frameworks such as perturbation cascade inference (PCI)  (?) and dynamical causal effect (DCE)  (?, ?) provide relevant computational schemes. To detect asymmetry information transfer in known two-dimensional dynamical systems, Liang and Kleeman proposed an analytic approach named Liang-Kleeman information flow  (?, ?, ?), which serves as a prototype for IntDC. This method quantifies causality by freezing one variable as a parameter, and evaluating the resulting outcomes using the Frobenius-Perron operator. The Liang information flow has been further extended to multivariate, stochastic, and quantum systems  (?, ?, ?, ?). However, due to ethical or practical limitations, many real systems should be analyzed without any external intervention. An essential problem is how to measure IntDC solely from the observational data but at the interventional level.
@@
In this study, we propose the IntDC framework and introduce a criterion named Interventional Embedding Entropy (IEE), which aims to identify and quantify IntDC between variables solely from the observational time series. Actually, IEE is rigorously derived by the theory of the delay embedding. Numerical IEE does not require specific prior knowledge of dynamics, and additional perturbation to the system is also unnecessary. Compared to ConDC indices, such as GC, TE, and CCM, IEE designed for IntDC has the capability to rank the importance of causal effects and construct directional causal networks more effectively. We demonstrate numerical experiments on both simulated examples and real datasets, including estimating neural connectomes of C. elegans, evaluating COVID-19 transmission in Japan, and constructing regulatory networks surrounding key circadian genes.
2 Methods
2.1 Constructive dynamical causality
The definition of ConDC for a complex system in the original time-series space is given as follows:
Definition 1 (Constructive dynamical causality, ConDC  (?)).
For a complex dynamical system
$\bm{x}_{t+1}=\bm{f}(\bm{x}_{t},\bm{x}_{t-1},\dots,\bm{x}_{t-p})+\bm{\varepsilon}_{t},$ (1)
where $\bm{x}=(x^{(1)},x^{(2)},\dots,x^{(n)})^{T}$ represents the system with $n$ components, $\bm{f}=(f_{1},f_{2},\dots,f_{n})^{T}$ is a vector function, $p$ is referred to the memory time, and $\bm{\varepsilon}_{t}$ is an independent noise term, there exists ConDC from component $x^{(j)}$ to $x^{(i)}$, if $\exists k\in\{1,2,\dots,p\}$ such that $\partial f_{i}/\partial x^{(j)}_{t-k}\neq 0$ for almost any $t$.
For simplicity, we consider a discrete two-variable dynamical system as an example:
@@ 2.1 Constructive dynamical causality
where $x,y$ are two variables, $p$ denotes the time step during which causality is considered, and $\varepsilon_{\cdot,t}$ stand for small noise terms. According to Definition 1, there exists ConDC from $x$ to $y$ since the evolving equation of $y_{t+1}$ depends on the historical behavior of $x$, while there is no ConDC from $y$ to $x$ as the dynamics of $x_{t+1}$ is independent of $y$. GC and TE measure the predictability of $f$ in the original time-series space to quantify the ConDC from $x$ to $y$ through linear regression and entropy uncertainty, respectively (Fig. 1(B)). The causal strength of GC/TE is based on the assumption of separability, which means that removing $x$ from the system does not influence the observed data of $y$. However, in general coupled systems, the separability is not satisfied, as removing one variable changes the values of affected ones  (?, ?, ?).
To address the challenge of detecting ConDC in universal non-separable systems, various approaches from the delay-embedding space have been proposed. In the delay-embedding framework, the bivariate system Eq. (2) is assumed to evolve into an attractive manifold with an inner dimension $d$. Let the time-delayed vectors of $x$ and $y$ be
@@
where $L$ is the time-delayed length, $\mathcal{M}_{X}$ and $\mathcal{M}_{Y}$ represent the manifolds formed by $\bm{X}_{t}$ and $\bm{Y}_{t}$, respectively. According to the seminal stochastic version of Takens’ embedding theorem  (?, ?, ?, ?), we can obtain the following theorem:
Theorem 1 (ConDC in delay-embedding space).
If $x$ is the ConDC of $y$ in dynamics Eq. (2), and $\bm{X}_{t}$, $\bm{Y}_{t}$ are the time-delayed vectors, respectively, then there exists a smooth projection operator $\bm{F}$ in generic sense such that
@@
when the time-delayed length satisfies $L\geqslant 2d$, where $d$ is the inner dimension of the attractive manifold.
@@
The detailed derivation can be referred to the Supplementary Text. The “generic sense” means that a smooth projection exists for a dense and open set of all possible time-delayed ways (?, ?). Equation (5) shows that the causal variable $\bm{X}_{t}\in\mathcal{M}_{X}$ can be reconstructed by the effect variable $\bm{Y}_{t+1}\in\mathcal{M}_{Y}$. However, as $\bm{F}$ is not reversible in generic sense, we can not determine $\bm{Y}_{t+1}$ only by the information from $\bm{X}_{t}$. According to Theorem 1, the causal dependence in $f$ between $x$ and $y$ is transformed into the reconstructability of $\bm{F}$ (Fig. 1(D)). Theorem1, established for the dynamical system in Eq. (2), can be generalized to multi-variable systems as described in Eq. (1), allowing for feedback interactions between different variables. Instead of fitting a model by removing the causal variable as in GC and TE, detecting the existence and quantifying the continuity characteristics of $\bm{F}$ in the delay-embedding space can be adequate for causal identification, especially for universal non-separable systems. Related algorithms include CCM, PCM, and EE.
@@
Figures 1(A), (B) and (D) summarize dynamical causality at the constructive level, where the ConDC from variable $x$ to $y$ is estimated only from the historical observed time series. But fitting Eq. (2) or Eq. (5) from data is criticized as mere association or prediction  (?, ?).
@@
Figure 1: Illustration for the constructive dynamical causality (ConDC) and interventional dynamical causality (IntDC). (A) An illustrative example of a complex system where $(x_{t},y_{t})$ are observed time series. The $z_{t}$ stands for the other variable that is not the primary focus of analysis. (B) ConDC from $x$ to $y$ indicates that the dynamics of $y$ depends on $x$. Algorithms such as Granger causality (GC) and transfer entropy (TE) infer the ConDC in the original time-series space. (C) IntDC from $x$ to $y$ indicates that the deviation of $y$, i.e. $\delta y$, depends on an intervention on $x$, i.e. $\delta x$. Usually additional data from a perturbed system $(\tilde{x}_{t},\tilde{y}_{t})$ is necessary for detecting IntDC in the original time-series space. The parameter $p$ is the time lag after intervention occurs. (D) ConDC can be detected in the embedding space by cross mapping $\mathbb{F}$. The causal vector $\bm{X}_{t}$ can be reconstructed by the effect vector $\bm{Y}_{t+1}$. Parameter $L$ is time delay used in the delay-embedding. (E) IntDC can be modeled in the embedding space. The interventional causal vector $\delta\bm{X}_{t}$ should be reconstructable by the effect vector $\delta\bm{Y}_{t+1}$ around $\bm{Y}_{t+1}$. In numerical computation, $\delta\bm{Y}_{t+1}$ is approximated by the distance from $\bm{Y}_{t+1}$ to its neighbors $\bm{Y}_{t_{k}+1}$, while $\delta\bm{X}_{t}$ is determined using $\bm{X}_{t_{k}}$ corresponding to the same time index $t_{k}$. (F) Interventional embedding entropy (IEE) measures the average information retention contained in $\delta\bm{Y}_{t+1}$ from $\delta\bm{X}_{t}$, quantifying the IntDC from variable $x$ to $y$. Notably, IEE does not require additional perturbations to the system and can be inferred from observational data solely.
2.2 Interventional dynamical causality
The detection of hidden essential-level causality necessitates intervention or manipulation to the system  (?, ?, ?, ?). In this study, we propose the definition of IntDC for a complex system in the original time-series space as the following:
Definition 2 (Interventional dynamical causality, IntDC).
For a complex dynamical system Eq. (1), one component $x^{(j)}$ is the IntDC of another component $x^{(i)}$, if $\exists k\in\{1,2,\dots,p\}$ such that $\delta x^{(i)}_{t+1}=\tilde{x}^{(i)}_{t+1}-x^{(i)}_{t+1}$ depends on $\delta x^{(j)}_{t-k}=\tilde{x}^{(j)}_{t-k}-x^{(j)}_{t-k}$ for almost any $t$, where $\tilde{x}^{(j)}_{t-k}$ is the intervened dynamics of $x^{(j)}$, and $\tilde{x}^{(i)}_{t+1}$ is the dynamics of affected $x^{(i)}$ after the intervention on $x^{(j)}$.
The intervention $\delta x^{(j)}_{t-k}$ can be either a pulse stimulation or persistent disturbances to the system.
Fig. 1(C) illustrates the IntDC in the time-series space for the simplified bivariate system Eq. (2), involving variables $x$ and $y$. To measure the IntDC, data before and after the intervention, i.e. $(x_{t},y_{t})$ and $(\tilde{x}_{t},\tilde{y}_{t})$, are typically required, as in PCI and DCE. PCI measures the distances of every node in the network from each perturbed node to reconstruct the directed causal network  (?). DCE classified causal coupling into four situations based on the type of effects (stationary statistic change, phase orbit change) and the type of interventions (state space interventions, parametric interventions)  (?). These conventional algorithms rely on reproducible dynamics or data availability under different settings.
@@ 2.2 Interventional dynamical causality
To address challenges of causal inference in widespread non-intervention, non-linear and non-separable systems, we turn to the study of intervened dynamics of Eq. (5) in the delay-embedding space. Denote $\delta\bm{X}_{t}=\widetilde{\bm{X}}_{t}-\bm{X}_{t}$ and $\delta\bm{Y}_{t+1}=\widetilde{\bm{Y}}_{t+1}-\bm{Y}_{t+1}$, where $\widetilde{\bm{X}}_{t}$ and $\widetilde{\bm{Y}}_{t+1}$ are the time-delayed vectors of $x$ and $y$ after the intervention, respectively. We can obtain the following theorem:
Theorem 2 (IntDC in delay-embedding space).
If $x$ is the IntDC of $y$ in dynamics Eq. (2), and $\bm{X}_{t},\bm{Y}_{t}$ are the time-delayed vectors, respectively, then there exists a smooth projection operator $\bm{F}$ in generic sense such that
$\delta\bm{X}_{t}=\nabla_{\bm{Y}}\bm{F}(\bm{Y}_{t+1})\cdot\delta\bm{Y}_{t+1},$ (6)
when the time-delayed length satisfies $L\geq 2d$ and the intervention is sufficiently small, where $d$ is the inner dimension of the attractive manifold.
Theorem 2 is directly deduced by Theorem 1. However, Eq. (6) provides insights on the causality under intervention. It indicates that $\delta\bm{Y}_{t+1}$ contains complete information of $\delta\bm{X}_{t}$ in the neighborhood of $\bm{Y}_{t+1}$ and can reconstruct $\delta\bm{X}_{t}$, if there exists IntDC from the causal variable $x$ to the effect variable $y$ (Fig. 1(E)). The conclusion in Eq. (6) of Theorem 2, derived from the dynamics in Eq. (2), remains valid for any pair of components exhibiting IntDC in the multi-variable system of Eq. (1), and feedback between different components is permitted. Figures 1(A), (C), and (E) give an overview of the IntDC framework.
2.3 Interventional embedding entropy